In July 2007, Apple had cash of $7.08 billion, current assets of $18.74 billion, current liabilities of $6.98 billion, and inventories of $0.21 billion.

(a) What was Apple's current ratio? Answer


(b) What was Apple's quick ration? Answer


(c) In July 2007, Dell had a quick ratio of 1.29 and a current ratio of 1.35. What can you say about the asset liquidity of Apple relative to Dell?

1. Apple has significantly more debt than Dell.

2. Dell has significantly more liquid assets than Apple relative to current liabilities.

3. Apple is more profitable than Dell.

4. Apple has significantly more liquid assets than Dell relative to current liabilities.
Answer #1

Hi

a) Current Ratio = Current asset/ current liabilities

= 18.74/6.98 = 2.69 times

b) Quick ratio = (Current Asset - inventories)/ Current liabilities

= (18.74- 0.21)/6.98 = 2.66 times

c) Since Apple Quick Ratio and current ratio is higher than Dell.

That means Apple has more liquid asset than Dell relative to current liabilities.

Hence option 4 is correct. Apple has significantly more liquid assets than Dell relative to current liabilities.

Thanks